# Construction of 2-D Z-Complementary Array Code Sets with Flexible Lengths for Different System Requirements

### v3

@@ Abstract
In this paper, we propose a new and optimal construction of two-dimensional (2-D) Z-complementary array code set (ZCACS) using multivariable extended Boolean functions (EBFs). The proposed 2-D arrays can be used in multi-carrier code division multiple access (MC-CDMA) systems. The proposed construction produces a better PMEPR upper bound than the existing constructions for such use cases. We also propose a tighter upper bound for the set size and showed that the proposed code set is optimal in special cases. Finally, We derive 2-D Golay complementary array set (GCAS) and Golay complementary set (GCS) from the proposed construction, which has application in uniform rectangular array (URA)-based massive multiple-input multiple-output (mMIMO) system to achieve omnidirectional transmission. The simulation result shows the performance benefits of the derived codes. In essence, we show that the flexibility of the parameters of the proposed 2-D ZCACS makes it a good candidate for a wide range of use cases in wireless communication.
Index Terms:
@@
Z-complementary array code set (ZCACS), peak-to-mean envelope power ratio (PMEPR), multi-carrier code division multiple access (MC-CDMA), multiple-input multiple-output (MIMO), uniform rectangular array (URA), space-time block code (STBC), omnidirectional transmission.
I Introduction
@@
The area of sequences design with good correlation properties for applications in wireless communication started with the discovery of the Golay complementary pairs (GCP), introduced in the seminal paper of Marcel J. E. Golay [1]. Leveraging the ideal auto-correlation sum spectrum of these sequences, researchers have found several applications of them in communication systems, for example, channel estimation [2], radar [3], reduction of peak-to-mean envelope power ratio (PMEPR) in orthogonal frequency division multiplexing (OFDM) [4, 5], etc. However, there are practical limitations on such sequences. For example, computer search shows that binary GCPs exist only for the lengths of the form $2^{\alpha}10^{\beta}26^{\gamma}$. To generalize GCP, Tseng et al. introduced the Golay complementary set (GCS) in 1972 [6], which has increased flexibility in sequence lengths. In the OFDM system, the length of the sequence determines the number of orthogonal sub-carriers. Nevertheless, this flexibility of lengths was achieved with the trade-off in an increased PMEPR, which is upper bounded by the number of sequences. A different direction of generalization was the introduction of zero-correlation zone (ZCZ) by Fan et al. [7], where the auto-correlation sum is zero in a specific zone of time-shifts. However, in a multiple-access system, an entire sequence set (also called code) is assigned to a user, and hence large numbers of codes are required which are orthogonal with respect to the correlation function. The research in this direction aims to achieve the optimality of the set size. For this purpose, mutually orthogonal Golay complementary set (MOGCS) [8], Z-complementary code set (ZCCS) [7], low-correlation zone (LCZ) sequence set [9], and quasi-complementary code set (QCCS)[10] were introduced.
@@
Recently, two-dimensional (2-D) array sets and code sets with their 2-D correlation properties have attracted significant attention due to their practical application benefits, such as multi-carrier code division multiple access (MC-CDMA). The traditional one-dimensional (1-D) correlation functions are special cases of 2-D correlation functions by definition. Hence, it is possible to use 2-D arrays in place of 1-D sequences. In such cases, the primary issue is power control, which depends on the correlation property. The 2-D correlation property of 2-D array sets enables us to control PMEPR in one dimension and achieve interference-free transmission in the other dimension. Also, Turcsány et al. in [11, 12] showed that relaxing the orthogonality property of a 2-D MOGCS, 2-D code sets with larger set size can be used in 2D-MC-CDMA. The benefit of using 2-D arrays instead of 1-D sequences in such a system is the increased freedom to position the array elements in the time and frequency dimensions simultaneously.
@@
Apart from it, some recent applications of 2-D array are the use of precoders for massive multiple-input multiple-output (mMIMO) to achieve omnidirectional transmission [13], ultra-wide band (UWB) communication [14, 15], etc. mMIMO has been extensively used in the 5th generation (5G) communication. It has the capability to increase the data rate by a large factor than the single-input single-output (SISO) system, making it an inseparable feature of future communication standards [16]. Hence, the literature has many works on 2-D code sets, such as 2-D Golay complementary array sets (GCAS) and 2-D quasi-complementary array sets (QCAS), for designing precoders in the mMIMO systems. The corresponding articles can be found here [17, 18, 19, 20, 21, 22]. However, applying 2-D QCAS as precoders has its limitations despite the increased flexibility of parameters. Because, the power radiation pattern is not omnidirectional here [22]. All this suggests that the existing 2-D codes does not completely solve the question of parameter flexibility in omnidirectional transmission. Another important parameter in such a system is the size of space-time block code (STBC) for encoding data [17]. Orthogonal STBCs with lower sizes are suitable for reducing complexity. It is an extremely difficult task to accommodate both MIMO and STBC configurations in a single construction of codes, due to the trade-off between the parameters. The best solution appears to be finding codes according to the requirements.
@@
In this paper, as the main contribution, we propose 2-D Z-complementary array code sets (ZCACS) of array size $\left(2^{m}p\right)\times\left(2p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}\right)$, where $m_{\alpha}\geq 2,\forall\alpha$, $p$, $p_{\alpha}$’s are prime numbers and $m\geq 1$. To do this, we first propose a construction of inter-group complementary (IGC) code set, which provides the necessary EBFs for the main construction. Then we analyze the PMEPR bounds and show that they are tighter than the existing constructions. We also investigate the set size upper bound and show that the proposed code set is optimal with respect to the bound for $p=2$. In addition to that, we derive suitable precoding matrices from the 2-D ZCACs for use in mMIMO to achieve omnidirectional transmission. Simulations are done to show the performance benefit of the precoders and parameter flexibility in URA and STBC.
@@
The rest of the paper is organized as follows: in Section II, the preliminary definitions and lemmas are discussed. In Section III, the main construction is proposed. In Section IV, the bound for PMEPR is analyzed. In Section V, the set size bound of 2-D ZCACS is investigated. In Section VI, the application benefit of 2-D code sets in mMIMO is discussed in detail. In Section VII, we compare the proposed construction with previous works. Finally, in Section VIII, we draw conclusion.
II Preliminaries
@@
In this section, we provide the definitions of the sequence/2-D array sets and code sets related to the work. Also, we give a structure of the EBFs used in the construction.
II-A Sequences and 2-D Arrays
@@
The sequences and 2-D arrays considered here are assumed to be unimodular, polyphase and complex-valued, which are used in wireless systems with $q$-ary phase shift keying ($q$-PSK) modulation. A 2-D array $\mathbf{X}$ of size $L_{1}\times L_{2}$ is denoted by $X_{i,j}$, where $0\leq i<L_{1}$, $0\leq j<L_{2}$. For $L_{2}=1$, it becomes a sequence of length $L_{1}$, denoted by a lowercase bold letter.
Definition 1 (2-D ACCF )
@@
$C(\mathbf{X},\mathbf{Y})(\tau_{1},\tau_{2})=\begin{cases}&\sum\limits_{i=0}^{L_{1}-1}\sum\limits_{j=0}^{L_{2}-1}X_{i,j}Y^{*}_{i+\tau_{1},j+\tau_{2}},\\ \end{cases}$ (1)
@@
Definition 2 (2-D Array Code)
@@
$\sum_{\eta=0}^{M-1}A(\mathbf{X}_{\eta})(\tau_{1},\tau_{2})=\begin{cases}ML_{1}L_{2},&(\tau_{1},\tau_{2})=(0,0);\\ 0,&0\leq\left|\tau_{1}\right|<Z_{1},\\ &0\leq\left|\tau_{2}\right|<Z_{2},\\ &(\tau_{1},\tau_{2})\neq(0,0).\end{cases}$ (3)
@@
Definition 5 (2-D ZCACS ([23]))
A collection of 2-D ZCAC $\mathcal{Z}=\{\mathcal{X}_{0},\mathcal{X}_{1},\dots,\mathcal{X}_{K-1}\}$ where $\mathcal{X}_{\nu}=\{\mathbf{X}_{0}^{\nu},\mathbf{X}_{1}^{\nu},\dots,\mathbf{X}_{M-1}^{\nu}\}$ and each $\mathbf{X}_{i}^{\nu}$ has equal size of $L_{1}\times L_{2}$, is called a 2-D ZCACS if they satisfy
$\begin{split}&\sum_{\eta=0}^{M-1}C(\mathbf{X}_{\eta}^{\nu_{1}},\mathbf{X}_{\eta}^{\nu_{2}})(\tau_{1},\tau_{2})\\ &=\begin{cases}ML_{1}L_{2},&\nu_{1}=\nu_{2},(\tau_{1},\tau_{2})=(0,0);\\ 0,&\nu_{1}=\nu_{2},(\tau_{1},\tau_{2})\neq(0,0),\\ &0\leq\left|\tau_{1}\right|<Z_{1},0\leq\left|\tau_{2}\right|<Z_{2};\\ 0,&\nu_{1}\neq\nu_{2},\left|\tau_{1}\right|<Z_{1},\left|\tau_{2}\right|<Z_{2};\\ \end{cases}\end{split}$ (4)
@@
Remark 1
1-D correlation functions and sequence sets appear as a special case of this 2-D versions. The corresponding definitions follow analogously with the change in $L_{1}=L,L_{2}=1$ and $Z_{1}=Z,Z_{2}=1$. In 1-D case, 2-D GCAS, GCAP, ZCAC, ZCAP, ZCACS are called Golay complementary set (GCS), GCP, Z-complementary set (ZCS), Z-complementary pair (ZCP), and Z-complementary code set (ZCCS), respectively.
Lemma 2 ([24, 25])
@@
Now, we define some special sequence sets which have been used in the construction process of 2-D ZCACS. We define those sequence sets in association with the 2-D counterparts as discussed in Remark 1.
Definition 6 (Complementary Mate)
@@
$\begin{split}\sum_{\eta=0}^{M-1}C(\mathbf{a}_{\eta}^{\nu_{1}},\mathbf{a}_{\eta}^{\nu_{2}})(\tau)=\begin{cases}ML,&\nu_{1}=\nu_{2},\tau=0;\\ 0,&\nu_{1}=\nu_{2},0<\left|\tau\right|<Z;\\ 0,&\nu_{1}\neq\nu_{2},\mathcal{C}_{\nu_{1}},\mathcal{C}_{\nu_{2}}\in\mathcal{S}^{g},\\ &\left|\tau\right|<Z;\\ 0,&\text{otherwise}.\end{cases}\\ \end{split}$ (8)
II-B Extended Boolean Function
Let $\mathbb{Z}_{n}=\{[0],[1],\dots,[n-1]\}$ be the ring of integers modulo $n$, where []delimited-[][$\cdot$][ ⋅ ] denote the residue class. Then any well-defined function $g:\mathbb{Z}_{L}\rightarrow\mathbb{Z}_{q}$ produces a sequence of length $L$, namely, $\left(g([0]),g([1]),\dots,g([L-1])\right)$ and it is the fundamental idea behind the sequence generation method using EBFs. Note that, any length $L$ can be written as $L=p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}$, where $p_{\alpha}$’s are distinct primes. Then the Chinese Remainder Theorem (CRT) says that $\exists$ a ring isomorphism $\sigma$ such that
@@ II-B Extended Boolean Function
given by $x\mapsto\left(x\mod p_{1}^{m_{1}},x\mod p_{2}^{m_{2}},\dots,x\mod p_{k}^{m_{k}}\right)$. Now, for the purpose of maintaining the well-defined property of EBFs, we assume $\mathbb{Z}_{n}$ as the set $\{0,1,\dots,n-1\}$ of numbers only, dropping the residue-class property. The corresponding functions rules are set accordingly to map the elements of $\mathbb{Z}_{L}$ to $\mathbb{Z}_{q}$, which we show shortly.
@@
For the sake of simplicity, we denote both types of extensions as $\mathbb{Q}_{q}$ and take either extension. It does not matter in the actual construction process.
@@
where $i$ is varied in the following manner: First $\vec{i}_{1}$ is varied from 00 to ${p^{\prime}}_{1}^{m_{1}}-1$, keeping others fixed at zero-vector. Then both $\vec{i}_{2}$ and $\vec{i}_{1}$ are varied, keeping others fixed. In summary, at $l$-th step $\{\vec{i}_{l},\dots,\vec{i}_{1}\}$ are varied, keeping the rest fixed, for $l=1,2,\dots,k$. This gives an enumeration of the integers $0\leq i<{p^{\prime}}_{1}^{m_{1}}{p^{\prime}}_{2}^{m_{2}}\dots p^{\prime}_{k}{}^{m_{k}}$ similar to the Boolean table.
@@
II-C 2-D Extended Boolean Function
@@
Similar to the 1-D case, we can also associate a complex-valued matrix $\Psi(K)$ with $K$, where $\left(\Psi(K)\right)_{i,j}=\omega_{q}^{K_{i,j}}$ for all $i,j$.
TABLE I: 2-D array corresponding to the functions $F_{\vec{d}}$
$K$= $\vec{d}=(d_{2},d_{3})$ $(v_{2,1},v_{3,1}$) (0,0) (1,0) (0,1) (1,1) (0,2) (1,2) (0,0) 0 0 0 0 0 0 (1,0) 0 2 0 2 0 2 (0,1) 0 0 3 3 2 2 (1,1) 0 2 3 1 2 0 (0,2) 0 0 2 2 0 0 (1,2) 0 2 2 0 0 2
Example 1
Let $F_{\vec{d}}:\mathbb{Z}_{2}\times\mathbb{Z}_{3}\rightarrow\mathbb{Q}_{4}$ be defined by $F_{\vec{d}}(v_{2,1},v_{3,1})=2v_{2,1}d_{2}+3v_{3,1}d_{3}$, where $\vec{d}=(d_{2},d_{3})$. Then the $6\times 6$ 2-D array $K$ related to $F_{\vec{d}}$ is generated in TABLE I.
II-D Peak-to-Mean Envelope Power Ratio
We consider MC-CDMA system, which combines the principles of both CDMA and OFDM. In an asynchronous system, column sequence of 1-D code is considered for power control [27]. For a 2-D code, a 2-D array of that code can also be used, using any one of the dimensions for power control. Given a $\mathbb{Q}_{q}$-valued sequence $\mathbf{a}=\{a_{0},a_{1},\dots,a_{L-1}\}$, the transmitted signal can be modeled as the real part of the complex envelope $S_{\mathbf{a}}(t)=\sum_{i=0}^{L-1}\omega_{q}^{a_{i}+qf_{i}t}$, where $f_{i}=f+i\Delta f$, $f$ is a constant, $\Delta f$ is an integer multiple of symbol rate and $0\leq\Delta ft\leq 1$. The term $\frac{\mid S_{\mathbf{a}}(t)\mid^{2}}{L}$ is called the instantaneous-to-average power ratio (IAPR). The PMEPR of the sequence $\mathbf{a}$ is given by

### v4

@@ Abstract
In this paper, we propose a new and optimal construction of two-dimensional (2-D) Z-complementary array code set (ZCACS) using multivariable extended Boolean functions (EBFs). The proposed 2-D arrays have many applications in modern wireless communications, such as multi-carrier code division multiple access (MC-CDMA), massive multiple input multiple output (mMIMO), etc. The main theoretical problem for sequences and 2-D arrays for application in MC-CDMA lies in the efficient construction of such sequences and arrays, which have low peak-to-mean envelope power ratio (PMEPR) and flexible parameter values. The PMEPR measures the power efficiency of the concerned system and hence has been an important research topic for past several years. The proposed construction produces a better PMEPR upper bound than the existing constructions. We also propose a tighter upper bound for the set size which translates more number of supported users in the communication system. We show that for some special cases, the proposed code set is optimal with respect to that bound. Finally, We derive 2-D Golay complementary array set (GCAS) and Golay complementary set (GCS) from the proposed construction, which has significant application in uniform rectangular array (URA)-based massive multiple-input multiple-output (mMIMO) system to achieve omnidirectional transmission. The simulation result shows the performance benefits of the derived arrays. In essence, we show that the flexibility of the parameters of the proposed 2-D ZCACS makes it a good candidate for practical use cases, both in theory and simulation.
keywords:
@@
Z-complementary array code set (ZCACS), extended Boolean function (EBF), peak-to-mean envelope power ratio (PMEPR), multi-carrier code division multiple access (MC-CDMA), multiple-input multiple-output (MIMO), uniform rectangular array (URA), space-time block code (STBC)
pacs:
MSC Classification]94A05, 94A15, 94A55
1 Introduction
@@
The area of sequence design with good correlation properties for applications in wireless communication started with the discovery of Golay complementary pairs (GCP), introduced in the seminal paper by Marcel J. E. Golay [GolayInfra]. Leveraging the ideal auto-correlation sum spectrum of these sequences, researchers have found several applications of them in communication systems, for example, channel estimation [ChannelEstm], radar [DopplerRadar], reduction of peak-to-mean envelope power ratio (PMEPR) in orthogonal frequency division multiplexing (OFDM) [Popo, DavisJedwab], etc. However, there are practical limitations on such sequences. For example, a computer search shows that binary GCPs exist only for lengths of the form $2^{\alpha}10^{\beta}26^{\gamma}$. To generalize GCP, Tseng et al. introduced the Golay complementary set (GCS) in 1972 [TsengCS], which has increased flexibility in sequence lengths. In the OFDM system, the length of the sequence determines the number of orthogonal subcarriers, and hence is an important parameter. Nevertheless, this flexibility of lengths was achieved with the trade-off in an increased PMEPR, which is upper-bounded by the number of constituent sequences.
A different direction of generalization was the introduction of zero-correlation zone (ZCZ) by Fan et al. [FanZCP], where the auto-correlation sum is zero in a specific zone around the zero time-shift. However, in a multiple-access system such as multi-carrier code division multiple access (MC-CDMA), an entire sequence set (also sometimes called a code) is assigned to a single user, and hence to support large number of users a large number of such codes are required, which are orthogonal with respect to the correlation function. Research in this direction aims to achieve the optimality of the set size. For this purpose, mutually orthogonal Golay complementary set (MOGCS) [hatoriCCC], Z-complementary code set (ZCCS) [FanZCP], low-correlation zone (LCZ) sequence set [LiuLCZ], and quasi-complementary code set (QCCS)[LiuQCSS] were introduced.
@@
Recently, two-dimensional (2-D) array sets and code sets with their 2-D correlation properties have attracted significant attention due to their practical application benefits, such as ultra wideband (UWB) [Zhang2DMOGCS, XengZCACS], 2D-MC-CDMA, massive multiple input multiple output (MIMO), etc. The traditional one-dimensional (1-D) correlation functions are special cases of 2-D correlation functions by definition. Hence, it is possible to use 2-D arrays in place of 1-D sequences in multi-carrier systems. The 2-D correlation property of 2-D array sets enables us to control PMEPR in one dimension and achieve interference-free transmission in the other dimension. In such cases, the primary issue is power control, which depends on the correlation property. Turcsány et al. in [Turcsny2004New2T, turcsanyQCAS] showed that relaxing the orthogonality property of a 2-D MOGCS, 2-D code sets with larger set size can be used in 2D-MC-CDMA. In such system, the benefit of using 2-D arrays instead of 1-D sequences is the increased freedom to position the array elements in the time and frequency dimensions simultaneously.
@@
Recently, applications of 2-D array as precoders for uniform rectangular array (URA)-based mMIMO [Omni1] has gained significant attention of researchers. The mMIMO system has been extensively used in the 5th generation (5G) of wireless communication. It has the capability to increase the data rate by a large factor in comparison to the single input single output (SISO) system, making it an inseparable feature of future communication standards as well [6Gsurvey]. The literature has many works on 2-D code sets for this application, such as 2-D Golay complementary array sets (GCAS) and 2-D quasi-complementary array sets (QCAS). The corresponding articles can be found here [OmniChen, YuboGCAS, YuboGCAS2, ZilongLiuOmni2024, YuboMIMO2024, YuboQCAS]. 2-D QCAS has an increased flexibility of parameter values in this regard, such as array size and flock size. However, applying 2-D QCAS as precoders has limitations. Because, the power radiation pattern is not omnidirectional in this case, which can be found here [YuboQCAS]. All this suggests that the existing 2-D codes does not completely solve the question of parameter flexibility in omnidirectional transmission. To achieve spatial diversity in such mMIMO system, space-time block code (STBC) is adopted [STBC, OmniChen], which is used to encode data before transmission. Orthogonal STBCs with lower sizes are suitable for reducing complexity. However, it is an extremely difficult task to achieve the desired flexibility of parameter values for both the URA and STBC in a single construction of codes, due to the underlying co-dependent relationship between the parameters values. The best solution appears to be finding codes according to the requirements.
@@
In this paper, as the main contribution, we propose 2-D Z-complementary array code sets (ZCACS) of array size $\left(2^{m}p\right)\times\left(2p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}\right)$, where $m_{\alpha}\geq 2,\forall\alpha$, $p$, $p_{\alpha}$’s are prime numbers and $m\geq 1$. To do this, we first propose a construction of inter-group complementary (IGC) code set, which provides the necessary EBFs for the main construction. Then we analyze the PMEPR bounds and show that they are tighter than the existing construction for similar array sizes. We also investigate the set size upper bound and prove a new upper bound for the set size of 2-D ZCACS. It can be seen that the proposed code set is optimal with respect to the bound for $p=2$. Considering the other aspect of application, i.e., mMIMO system, we derive suitable precoding matrices from the 2-D ZCACs to achieve omnidirectional transmission. Simulations are done to show the performance benefit of the precoders and parameter flexibility in URA and STBC.
@@
The rest of the paper is organized as follows: in Section 2, the preliminary definitions and necessary lemmas are discussed. In Section 3, the main construction of 2-D ZCACS is proposed. In Section 4, the bound for PMEPR is analyzed. In Section 5, the set size bound of 2-D ZCACS is investigated. In Section 6, the application benefit of 2-D code sets in mMIMO is discussed in detail. In Section 7, we compare the proposed construction with previous works with respect to the parameter values. Finally, in Section 8, conclusions are drawn.
2 Preliminaries
@@
In this section, we discuss some definitions of the sequence/2-D array sets and code sets related to the work. Also, we give a structure of the EBFs used in the construction.
2.1 Sequences and 2-D Arrays
@@
The sequences and 2-D arrays considered here are assumed to be unimodular, polyphase and complex-valued, which are used in wireless systems with $q$-ary phase shift keying ($q$-PSK) modulation. A 2-D array $\mathbf{X}$ of size $L_{1}\times L_{2}$ is denoted by $X_{i,j}$, where $0\leq i<L_{1}$, $0\leq j<L_{2}$. For $L_{2}=1$, it becomes a sequence of length $L_{1}$, usually denoted by a lowercase bold letter.
Definition 1 (2-D ACCF ).
@@
$C(\mathbf{X},\mathbf{Y})(\tau_{1},\tau_{2})=\sum\limits_{i=0}^{L_{1}-1}\sum\limits_{j=0}^{L_{2}-1}X_{i,j}Y^{*}_{i+\tau_{1},j+\tau_{2}},$ (1)
@@
Definition 2 (2-D Array Code).
@@
$\sum_{\eta=0}^{M-1}A(\mathbf{X}_{\eta})(\tau_{1},\tau_{2})=\begin{cases}ML_{1}L_{2},&(\tau_{1},\tau_{2})=(0,0);\\ 0,&0\leq\left|\tau_{1}\right|<Z_{1},0\leq\left|\tau_{2}\right|<Z_{2},\\ &(\tau_{1},\tau_{2})\neq(0,0).\end{cases}$ (3)
@@
Definition 5 (2-D ZCACS ([ZengZCACS2])).
A collection of 2-D ZCACs $\mathcal{Z}=\{\mathcal{X}_{0},\mathcal{X}_{1},\dots,\mathcal{X}_{K-1}\}$ where $\mathcal{X}_{\nu}=\{\mathbf{X}_{0}^{\nu},\mathbf{X}_{1}^{\nu},\dots,\mathbf{X}_{M-1}^{\nu}\}$ and each $\mathbf{X}_{i}^{\nu}$ has equal size of $L_{1}\times L_{2}$, is called a 2-D ZCACS if they satisfy
[…]
@@
Remark 1.
1-D correlation functions and sequence sets appear as a special case of these 2-D correlation functions and 2-D array sets. The corresponding definitions follow analogously only with the change of notations: $L_{1}=L,L_{2}=1$ and $Z_{1}=Z,Z_{2}=1$. The 1-D counterparts of 2-D GCAS, GCAP, ZCAC, ZCAP, ZCACS are called Golay complementary set (GCS), GCP, Z-complementary set (ZCS), Z-complementary pair (ZCP), and Z-complementary code set (ZCCS), respectively.
[…]
Lemma 2 ([FengZCCSbound]).
@@
Now, we define two special sequence sets which have been used in the construction of the proposed 2-D ZCACS.
Definition 6 (Complementary Mate).
@@
[…]
2.2 Extended Boolean Function
In this subsection, we introduce the mathematical tool that has been used for the proposed construction. Let $\mathbb{Z}_{n}=\{[0],[1],\dots,[n-1]\}$ be the ring of integers modulo $n$, where $\cdot$] denote the residue class. Then any well-defined function $g:\mathbb{Z}_{L}\rightarrow\mathbb{Z}_{q}$ produces a sequence of length $L$, namely, $\left(g([0]),g([1]),\dots,g([L-1])\right)$ and it is the fundamental idea behind the sequence generation method using EBFs.
Note that, any length $L$ can be written as $L=p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}$, where $p_{\alpha}$’s are distinct primes. Then the Chinese Remainder Theorem (CRT) says that $\exists$ a ring isomorphism $\sigma$ such that
@@ 2.2 Extended Boolean Function
given by $x\mapsto\left(x\mod p_{1}^{m_{1}},x\mod p_{2}^{m_{2}},\dots,x\mod p_{k}^{m_{k}}\right)$.
Now, for the purpose of maintaining the well-defined property of EBFs, we assume $\mathbb{Z}_{n}$ as the set $\{0,1,\dots,n-1\}$ of numbers only, dropping the residue-class property. The corresponding functions rules are set accordingly to map the elements of $\mathbb{Z}_{L}$ to $\mathbb{Z}_{q}$, which we show shortly.
@@
For the sake of simplicity, we denote both types of extensions as $\mathbb{Q}_{q}$ and take either extension. It does not matter in the actual construction process, as we shall see later.
@@
where $i$ is varied in the following manner: First $\vec{i}_{1}$ is varied from 0 to ${p^{\prime}}_{1}^{m_{1}}-1$, keeping others fixed at zero-vector. Then both $\vec{i}_{2}$ and $\vec{i}_{1}$ are varied, keeping others fixed. In summary, at $l$-th step $\{\vec{i}_{l},\dots,\vec{i}_{1}\}$ are varied, keeping the rest fixed, for $l=1,2,\dots,k$. This gives an enumeration of the integers $0\leq i<{p^{\prime}}_{1}^{m_{1}}{p^{\prime}}_{2}^{m_{2}}\dots p^{\prime}_{k}{}^{m_{k}}$ similar to a Boolean-representation table.
@@
2.3 2-D Extended Boolean Function
@@
Similar to the 1-D case, we can also associate a complex-valued matrix $\Psi(K)$ with $K$, where $\left(\Psi(K)\right)_{i,j}=\omega_{q}^{K_{i,j}}$ for all $i,j$. Below, we show an example of this for ease of understanding.
Example 1.
Let $F_{\vec{d}}:\mathbb{Z}_{2}\times\mathbb{Z}_{3}\rightarrow\mathbb{Q}_{4}$ be defined by $F_{\vec{d}}(v_{2,1},v_{3,1})=2v_{2,1}d_{2}+3v_{3,1}d_{3}$, where $\vec{d}=(d_{2},d_{3})$. Then the $6\times 6$ 2-D array $K$ related to $F_{\vec{d}}$ is generated in Table 1.
Table 1: 2-D array corresponding to the functions $F_{\vec{d}}$
$\vec{d}=(d_{2},d_{3})$ $(v_{2,1},v_{3,1}$) (0,0) (1,0) (0,1) (1,1) (0,2) (1,2)
(0,0) 0 0 0 0 0 0
(1,0) 0 2 0 2 0 2
(0,1) 0 0 3 3 2 2
(1,1) 0 2 3 1 2 0
(0,2) 0 0 2 2 0 0
(1,2) 0 2 2 0 0 2
2.4 Peak-to-Mean Envelope Power Ratio
In this subsection, we discuss the problem of PMEPR reduction in multi-carrier systems. In a MC-CDMA system, column sequence of 1-D code is considered for power control [LiuMCCDMA]. For a 2-D code, a 2-D array of that code can also be used, using any one of the dimensions for power control. Given a $\mathbb{Q}_{q}$-valued sequence $\mathbf{a}=(a_{0},a_{1},\dots,a_{L-1})$, the transmitted signal is modeled as the real part of the complex envelope
$S_{\mathbf{a}}(t)=\sum_{i=0}^{L-1}\omega_{q}^{a_{i}+qf_{i}t},$ (18)
where $_{i}=f+i\Delta f, ff$ is a constant, $\Delta f$ is an integer multiple of symbol rate and $0\leq\Delta ft\leq 1$ [DavisJedwab]. The term $\left|S_{\mathbf{a}}(t)\right|^{2}/L$ is called the instantaneous-to-average power ratio (IAPR). The PMEPR of the sequence $\mathbf{a}$ is given by